OpenStudy (anonymous):

what is the area of a triangle with a base of 7 in . and height of 12in. ? a.42 in2 b.19in2 c.84 in2 ? d.9.5 in2

OpenStudy (anonymous):

A=(1/2)bh = .5(7)(12)= 42 in^2

OpenStudy (anonymous):

triangle's area=1/2bh 1/2(7x12) 42

OpenStudy (anonymous):

42